71. Cox, North, and Lee form a partnership. Cox contributes $189,000, North contributes $157,500, and Lee contributes $283,500. Their partnership agreement calls for a 5% interest allowance on the partner's capital balances with the remaining income or loss to be allocated equally. If the partnership reports income of $184,500 for its first year, what amount of income is credited to North's capital account?

  • $61,500.

  • $65,175.

  • $60,450.

  • $51,000.

  • $58,875.

78. Mohr Company purchases a machine at the beginning of the year at a cost of $38,000. The machine is depreciated using the double-declining-balance method. The machine’s useful life is estimated to be 5 years with a $7,000 salvage value. Depreciation expense in year 2 is:

  • $7,600.

  • $9,120.

  • $15,200.

  • $22,800.

  • $12,400.

80. A company has bonds outstanding with a par value of $100,000. The unamortized premium on these bonds is $2,300. If the company retired these bonds at a call price of 98, the gain or loss on this retirement is:

  • $2,300 loss.

  • $2,000 gain.

  • $4,300 gain.

  • $2,300 gain.

  • $2,000 loss.

84. Bloom and Plant organize a partnership on January 1. Bloom's initial investment consists of $900 cash, $2,400 equipment and a $700 note payable reflecting a bank loan for the new business. Plant's initial investment is cash of $4,000. These amounts are the values agreed on by both partners. The journal entry to record Plant's investment is:

  • Debit Cash $4,000; credit Note Payable $700, credit Plant, Capital $3,300.

  • Debit Cash $3,300; debit Note Payable $700; credit Plant, Capital $4,000.

  • Debit Cash $4,700; credit Note Payable $700; credit Plant, Capital $4,700.

  • Debit Cash $4,000; credit Plant, Capital $4,000.

  • Debit Bloom, Capital $4,000; credit Cash $4,000
Answer #1

71) Calculation of Interest

Partners

Capital Balance

Rate of interest

Interest

Cox

$      189,000.00

5%

$    9,450.00

North

$      157,500.00

5%

$    7,875.00

Lee

$      283,500.00

5%

$ 14,175.00

Total

$      630,000.00

$ 31,500.00

Remaining Profit After Interest = 184,500 - 31,500 = 153,000

North Share = 153,000 x 1/3 = 51,000

Total Share of North = 7,875 +51,000 = 58,875

78) Double Declining Rate = 1/5*2 = 40%

First Year Depreciation Expense = 38,000 x 40% = 15,200

Second Year Depreciation Expense = (38,000 -15,200) x 40% = 9,120

80)Bonds Carrying Value = 100,000 +2,300 = 102,300

Repurchase Price = 100,000* 98% = 98,000

Gain = 102,300 - 98,000 = 4,300

84) Debit Cash $4,000; credit Plant, Capital $4,000
Answer #2

Let's calculate the amount of income credited to North's capital account and the depreciation expense for year 2 using the given information:

Question 1:Cox, North, and Lee form a partnership. Their partnership agreement calls for a 5% interest allowance on the partner's capital balances with the remaining income or loss to be allocated equally. The partnership reports income of $184,500 for its first year.

First, let's calculate the total capital contributed by all partners: Total capital = Cox's contribution + North's contribution + Lee's contribution Total capital = $189,000 + $157,500 + $283,500 Total capital = $630,000

Next, let's calculate the 5% interest allowance on each partner's capital: Interest allowance = 5% of each partner's capital Interest allowance for Cox = 5% of $189,000 = $9,450 Interest allowance for North = 5% of $157,500 = $7,875 Interest allowance for Lee = 5% of $283,500 = $14,175

Now, let's calculate the remaining income after deducting the interest allowances: Remaining income = Total income - (Interest allowance for Cox + Interest allowance for North + Interest allowance for Lee) Remaining income = $184,500 - ($9,450 + $7,875 + $14,175) Remaining income = $152,000

Since the remaining income is to be allocated equally, the amount of income credited to each partner's capital account is: Income credited to North's capital account = Remaining income / Number of partners Income credited to North's capital account = $152,000 / 3 Income credited to North's capital account = $50,666.666...

Rounded to two decimal places, the amount of income credited to North's capital account is approximately $50,666.67.

Answer:The amount of income credited to North's capital account is $50,666.67.

Question 2:Mohr Company purchases a machine at the beginning of the year at a cost of $38,000. The machine is depreciated using the double-declining-balance method. The machine’s useful life is estimated to be 5 years with a $7,000 salvage value. Depreciation expense in year 2 is:

First, let's calculate the depreciation rate: Depreciation rate = 2 / Useful life Depreciation rate = 2 / 5 Depreciation rate = 0.4

Now, let's calculate the depreciation expense for year 2 using the double-declining-balance method: Depreciation expense for year 1 = (Cost - Accumulated depreciation) * Depreciation rate Depreciation expense for year 1 = ($38,000 - $0) * 0.4 Depreciation expense for year 1 = $15,200

Depreciation expense for year 2 = (Cost - Accumulated depreciation) * Depreciation rate Depreciation expense for year 2 = ($38,000 - $15,200) * 0.4 Depreciation expense for year 2 = $9,120

Answer:The depreciation expense in year 2 is $9,120.

answered by: Hydra Master